# Optimal CO2 storage management considering safety constraints in multi-stakeholder multi-site CCS projects: a game theoretic perspective

### v1

@@ 6 Conclusions
A safe Multi-Agent Deep Deterministic Policy Gradient (MADDPG) framework based on a constrained Markov game has been developed to optimize each stakeholder’s objectives in geological carbon storage (GCS) operations while ensuring compliance with safety constraints. The proposed approach, although not strictly guaranteeing Pareto-optimality, consistently delivers strong performance in both competitive and cooperative scenarios.
In optimal basin-scale CCS management, reward and penalty designs play a central role in maximizing stakeholders’ objectives within safety constraints. In this work, reward models based on fully cooperative and fully competitive settings are discussed. Two penalty model strategies were investigated – point-wise and region-wide – and the results show that applying project-wide constraints is essential to prevent pressure trespassing from one operator’s lease area into another, particularly when operators have varying risk tolerances. This ensures a more equitable and safe operational environment. Overall, the safe MADDPG effectively balances economic objectives with safety requirements, demonstrating its potential as a robust decision-making tool for multi-stakeholder $CO_{2}$ storage projects with safety requirements.
In future work, we aim to investigate partially cooperative scenarios, where certain subsets of stakeholders collaborate while simultaneously competing with others. Such mixed cooperation–competition settings better reflect the complexities of real-world coalition agreements in CCS operations. In addition, we plan to develop multi-agent reinforcement learning algorithms with explicit Pareto-optimality guarantees, ensuring that learned strategies achieve efficient trade-offs where no stakeholder’s outcome can be improved without reducing others’. These advancements will enhance the framework’s applicability to more realistic operational contexts.
@@ Acknowledgments
We are grateful to the Gulf Coast Carbon Center at UT Austin and its sponsors for funding this work. We thank Sean Avitt for providing the Wilcox Group geomodel in Figure 5. Thanks also go to Computer Modeling Group Ltd. for providing the simulation software.
@@
Figure 15: Safe MADDPG with centralized training and decentralized execution (CTDE) (Lowe et al., 2017). Inside the Execution box, each agent $i$ selects $a_{i}=\pi_{i}(o_{i})$ from its local observation. During Training (outer box), a reward critic $Q^{r}_{i}$ and a cost critic $Q^{c}_{i}$ for each agent condition on the joint tuple $(o_{1:n},a_{1:n})$ and backpropagate gradients to the actors. Safety is enforced via a Lagrangian penalty $\lambda_{i}Q^{c}_{i}$, with multipliers $\lambda_{i}$ updated to penalize constraint violations; at evaluation time only the policies run, while critics (and $\lambda_{i}$ updates) are not required.
(a) Optimal well injection strategies over 25 years from safe MADDPG
(b) Optimal well injection strategies over 25 years from safe MADDPG
(c) Optimal well injection strategies over 25 years from CMOO
(d) Optimal well injection strategies over 25 years from CMOO
Figure 16: Optimal well injection strategies over 25 years from two approaches (a) fully collaborative scenario and (b) fully competitive scenario (c) fully collaborative scenario and (d) fully competitive scenario.
@@
7:   Execute joint action $\mathbf{a}=(a_{1},\ldots,a_{N})$; observe reward vector $\mathbf{r}$, penalty vector $\mathbf{c}$ and next observations $\mathbf{o}^{\prime}$ according to equations 4, 1619.

### v2

@@ 6 Conclusions
This study develops a constrained Markov game (CMG) formulation and a safe multi-agent deep deterministic policy gradient (MADDPG) implementation for basin-scale geological carbon storage (GCS), where multiple operators interact through shared pressure dynamics and must satisfy safety constraints. To make repeated forward evaluations tractable, we couple the learning framework with an Embed-to-Control (E2C) reduced-order surrogate, enabling efficient training and rollout in a sequential decision-making setting.
Using a three-operator synthetic basin case study, we show that coalition structure strongly shapes both system-level efficiency and the distribution of economic outcomes. Under full cooperation (grand coalition), agents share a team reward and learn coordinated controls that yield the highest total NPV in our experiments (\$12,191 M), whereas full competition produces the lowest total NPV (\$8,942 M), consistent with inefficiencies induced by self-interested responses under interference. In mixed coalition settings, partial cooperation systematically benefits the coalition members relative to the fully competitive baseline while disadvantaging the standalone operator, revealing persistent payoff asymmetries that can emerge when coordination is selective rather than basin-wide. Across all scenarios, the safety-aware learning formulation (cost critics with Lagrangian updates) leads to policies that respect the prescribed pressure limits, supporting the use of CMG/safe-MARL as a practical mechanism for learning decentralized, executable operating rules under regulatory constraints. We further benchmark CMG/safe-MADDPG against constrained multi-objective optimization (CMOO) using NSGA-II. The MOO baseline provides interpretable, preference-controlled trade-offs and yields comparatively balanced allocations, while the cooperative MADDPG outcome achieves the highest total NPV and is non-dominated within the discrete set of solutions reported.
Our future endeavor will focus on addressing some limitations of current study. First, while pressure is a widely used and illustrative safety constraint, the proposed CMG framework is not restricted to pressure and can incorporate additional operational or regulatory constraints (e.g., plume extent/AoR, leakage-risk indicators) through appropriate cost definitions.Second, the case study is designed to test the decision-making paradigm; scaling to field settings will require robust surrogates (or other efficient forward models), proper adjustment of the code algorithm, and so on. Finally, although the cooperative MADDPG solution is Pareto-efficient within the compared discrete set, establishing global Pareto optimality over the continuous feasible space would require explicit verification against the full Pareto front. Developing MARL variants with stronger Pareto-efficiency guarantees and/or integrating preference articulation into learning are promising directions.
@@ Acknowledgments
We acknowledge the Gulf Coast Carbon Center at UT Austin and its sponsors for funding this work. We thank Sean Avitt for providing the Wilcox Group geomodel in Figure 5. Thanks also go to Computer Modeling Group Ltd. for providing the simulation software. Finally, we appreciate the helpful discussions with Dr. Per Pettersson and Dr. Sarah Gasda regarding game-theoretic applications in CCS projects.
@@
Declaration of generative AI and AI-assisted technologies in the manuscript preparation process
During the preparation of this manuscript, the authors used ChatGPT 5.1 to assist with language editing. All content generated with this tool was subsequently reviewed, revised, and approved by the authors, who take full responsibility for the final published work.
@@
Figure 12 illustrates the graphical structure of the E2C model, where the gray boxes denote the training tuple $\{x_{t},\,u_{t},\,x_{t+1},\,y_{t+1}\}$. The encoder network $Q_{\varphi}$ maps inputs into a latent representation, while the decoder network $P_{\theta}$ reconstructs the corresponding outputs. The transition model, expressed as $F_{\omega_{1}}$, predicts the next latent state $z_{t+1}$ from the current latent state $z_{t}$ and the control input $u_{t}$. In practice, this transition function is implemented using two neural networks that approximate the matrices $A$ and $B$, thereby enabling the use of a locally linear dynamical equation. Likewise, the output network $G_{\omega_{2}}$ is designed to approximate the matrices $C$ and $D$. For further implementation and validation details, readers are referred to [13].
Figure 12: Graphical representation of E2C model, adapted from [13]
Figure 13: Safe MADDPG with centralized training and decentralized execution (CTDE) [35]. Inside the Execution box, each agent $i$ selects $a_{i}=\pi_{i}(o_{i})$ from its local observation. During Training (outer box), a reward critic $Q^{r}_{i}$ and a cost critic $Q^{c}_{i}$ for each agent condition on the joint tuple $(o_{1:n},a_{1:n})$ and backpropagate gradients to the actors. Safety is enforced via a Lagrangian penalty $\lambda_{i}Q^{c}_{i}$, with multipliers $\lambda_{i}$ updated to penalize constraint violations; at evaluation time only the policies run, while critics (and $\lambda_{i}$ updates) are not required.
Figure 13 shows the centralized training decentralized execution scheme in MADDPG. Figure 14 displays the well control strategies of all 3 companies over 20 years, while figure 15 presents the corresponding reservoir pressure distribution at the end of injection.
@@
7:   Execute joint action $\mathbf{a}=(a_{1},\ldots,a_{N})$; observe reward vector $\mathbf{r}$, penalty vector $\mathbf{c}$ and next observations $\mathbf{o}^{\prime}$ according to equations 6, 21 and 4, .